OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

OpenStudy (anonymous):

helppppp

OpenStudy (gorv):

find the derivative

OpenStudy (gorv):

\[\frac{ ds(t) }{ dt }=0-6=-6\]

OpenStudy (gorv):

so velocity is constant at t=2 it will be -6